show how effective atomic number explain the stability of complex compound?

Grade:Upto college level

2 Answers

Chetan Mandayam Nayakar
12 years ago

 

Stronger is the metal-ligand bond, less is the dissociation in the solution and hence greater is the stability of a coordination compounds.
Instability constant for the complex ion [Cu(NH3)4]2+ i.e.
[Cu(NH3)4]2+  ? Cu2+ + 4NH3, is given by the expression; Ki = [Cu2+][NH3]4/[Cu(NH3)4]2+
Stability constant of the above complex i.e.
Cu2+ + 4NH3 ? [Cu(NH3)2+ is given as under ; K = [Cu(NH3)42+]/[Cu2+][NH3]4 = 1/K,
Greater is the stability constant, stronger is the metal – ligand bond
Factors affecting the stability of complex ion
(1) Nature of central metal ion : The higher the charge density on the central metal ion the greater is the stability of the complex
For example, the stability constant of [Fe(CN)6]3– is much greater than the stability constant of [Fe(CN)6]4–.
         Fe2+ + 6CN ? [Fe(CN)6]4–; K = 1.8 × 106
         Fe3+ + 6CN ? [Fe(CN)6]3–; K = 1.2 × 1031

Effective atomic number (EAN) or Sidgwick theory: In order to explain the stability of the complexes sidgwick proposed effective atomic number. EAN generally coincides with the atomic number of next noble gas in some cases. EAN is calculated by the following relation :

EAN = Atomic no. of the metal –e lost in ion formation +No. of e gained from the donor atom of the ligands.

EAN = Atomic number – Oxidation number + co-ordination no. × 2
Prajnansh gulzar
5 years ago
Effective atomic number does not tell about the stability of compound it was jus suggested by sidgwick as it so, but further researches shows that there are lot of coordination compounds which contradict EAN rule thus such compounds become limitation to this rule like [Fe(CN)6]3- does not have nobel gas configuration but exist in nature. Nowadays EAN rule just tells about the electron density of central metal atom and reletive stabilty of two compunds even of those compounds in which one compound achieves EAN as the AN of nearest nobel gas otherwise the stability can be calculated only by the stability coefficient individually. Hence , EAN does not tell about wether a compound is stable or not.